# Itô’s formula for flows of conditional measures on semimartingales

### v2

@@ 3.3. Construction of conditional independent copies: general cases
and the extended $\sigma$-algebra $\overline{\mathcal{F}}$
¯$\overline{\mathcal{F}}=\sigma\{A_{0}\times A_{1}\times\cdots\times A_{n}:A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}\}.$ (3.7)
To ensure that these copies are conditional independent, let us define the probability measure $\overline{\mathbb{P}}$ as follows: for $A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}$, define
¯$\overline{\mathbb{P}}\big{(}A_{0}\times A_{1}\times\cdots\times A_{n}\big{)}:={\mathbb{E}}\bigg{[}\mathbbm{1}_{A_{0}}\prod_{i=1}^{n}{\mathbb{P}}(A_{i}|{\mathcal{G}})\bigg{]}.$ (3.8)
This measure is properly defined since the set $\{A_{0}\times A_{1}\times\cdots\times A_{n}\}$ is a $\pi$-system. Finally, define ~$\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$ for the extension of ${\mathcal{F}}$ and ${\mathcal{G}}$ into this enlarged probability space as
~:=$\widetilde{\mathcal{F}}:=\{A\times\Omega^{n}:A\in{\mathcal{F}}\},\quad\widetilde{\mathcal{G}}:=\{A\times\Omega^{n}:A\in{\mathcal{G}}\}.$ (3.9)
@@
Given a probability space $(\Omega,{\mathcal{F}},{\mathbb{P}})$ and a sub $\sigma$-algebra ${\mathcal{G}}\subset{\mathcal{F}}$. Then there exists uniquely an enlarged probability space $(\overline{\Omega},\overline{\mathcal{F}},{\mathbb{P}})$ satisfying (3.6), (3.7) and (3.8); and for any random variable $R$ in $(\Omega,{\mathbb{F}},{\mathbb{P}})$, there exist an $\widetilde{R}$ and $n$ conditional independent copies $\{R^{(i)}\}_{i=1,2,..n}$ by
@@
for $\overline{\omega}=(\omega_{0},,\omega_{1},\cdots,\omega_{n})\in\overline{\Omega}$ and $\widetilde{R}$, $\{R^{(i)}\}_{i=1,2,..n}$ satisfying
  • $\widetilde{R},\big{\{}R^{(i)}\big{\}}_{i=1,2,..n}$ are independent given $\widetilde{\mathcal{G}}$,
  • $\widetilde{R},\big{\{}R^{(i)}\big{\}}_{i=1,2,..n}$ are independent given $\widetilde{\mathcal{F}}$,
  • Law$\big{(}R\big{|}{\mathcal{G}}\big{)}(\omega_{0})=$Law$\big{(}\widetilde{R}\big{|}\widetilde{\mathcal{G}}\big{)}(\overline{\omega})=$Law$\big{(}R^{(i)}\big{|}{\widetilde{\mathcal{G}}}\big{)}(\overline{\omega})=$Law$\big{(}R^{(i)}\big{|}{\widetilde{\mathcal{F}}}\big{)}(\overline{\omega})$, for all $i=1,2,..,n$, a.s.
with $\overline{\omega}=(\omega_{0},\cdots,\omega_{n})\in\overline{\Omega}$ and $\widetilde{\mathcal{F}}$, $\widetilde{\mathcal{G}}$ defined by (3.9).
Note that when ${\mathcal{G}}$ is a trivial $\sigma$-algebra $\{\emptyset,\Omega\}$, then the above theorem reduces to the standard construction of independent copies. Its proof is given in Section 5.1.
Now, in the context of semimartingales, there is a corresponding version of the conditional independent copy where the probability space is a filtered space and sub $\sigma$-algebra ${\mathcal{G}}={\mathcal{G}}_{T}$ comes from the sub filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}$. Now define $\overline{\mathbb{F}}:=(\overline{\mathcal{F}}_{t})_{t\in(0,T)}$ by
¯$\overline{\mathcal{F}}_{t}=\sigma\{A_{0}\times A_{1}\times\cdots\times A_{n}:A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}_{t}\},$ (3.11)
and further define $\widetilde{\mathbb{F}}=(\widetilde{\mathcal{F}}_{t})_{t\in(0,T)}$ and $\widetilde{\mathbb{G}}=(\widetilde{\mathcal{G}}_{t})_{t\in(0,T)}$ respectively for the extension of ${\mathbb{F}}$ and ${\mathbb{G}}$ into this enlarged probability space as
~$\widetilde{\mathcal{F}}_{t}:=\{A\times\Omega^{n}:A\in{\mathcal{F}}_{t}\},\quad\widetilde{\mathcal{G}}_{t}:=\{A\times\Omega^{n}:A\in{\mathcal{G}}_{t}\}.$ (3.12)
@@
Given two ${\mathbb{R}}^{d}$-valued ${\mathbb{F}}$-semimartingales $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ and $\bm{Y}=(\bm{X}_{t})_{t\in[0,T]}$ on a filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$, and a sub $\sigma$-algebra ${\mathcal{G}}\subset{\mathcal{F}}$, there exists uniquely an enlarged probability space $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},\overline{\mathbb{P}})$ satisfying (3.6), (3.7), (3.8) and (3.11). Moreover, one can define two conditional independent copies $(\bm{X}^{\prime},\bm{Y}^{\prime})$, $(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})$ of the random processes $(\bm{X},\bm{Y})$ by
@@
and $({\bm{X}},{\bm{Y}})$, $(\bm{X}^{\prime},\bm{Y}^{\prime})$ and $(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})$ satisfying
$\begin{cases}\text{Law}\big{(}{\bm{X}},{\bm{Y}}\big{|}{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime},\bm{Y}^{\prime}\big{|}{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime},\bm{Y}^{\prime}\big{|}{\mathcal{F}}\big{)}\\ \qquad=\text{Law}\big{(}\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime}\big{|}{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime}\big{|}{\mathcal{F}}\big{)}\text{ a.s.}\\ ({\bm{X}},{\bm{Y}}),(\bm{X}^{\prime},\bm{Y}^{\prime}),(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})\text{ are independent given }{\mathcal{G}},\text{and are also independent given }{\mathcal{F}}.\end{cases}$ (3.13)
Furthermore, if the sub $\sigma$-algebra ${\mathcal{G}}={\mathcal{G}}_{T}$ from a sub filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$ satisfies the compatibility conditions as in Assumption 2.5, then this compatibility condition holds for the natural extension of $\overline{\mathbb{F}}$ and $\overline{\mathbb{G}}$ defined by (3.12) in the extended probability space. That is, for each $0\leq t\leq T$, ~$\widetilde{\mathcal{F}}_{t}$ and $\widetilde{\mathcal{G}}_{T}$ are independent given $\overline{\mathcal{G}}_{t}$, written as ~$\widetilde{\mathcal{F}}_{t}\perp\!\!\!\perp\widetilde{\mathcal{G}}_{T}\big{|}\widetilde{\mathcal{G}}_{t}.$
Note that Theorem 3.1 allows us to use the notation $(\bm{X},\bm{Y}),{\mathcal{F}}$, and ${\mathcal{G}}$ to respectively represent $(\widetilde{\bm{X}},\widetilde{\bm{Y}}),\widetilde{\mathcal{F}}$, and $\widetilde{\mathcal{G}}$ without any ambiguity, as is adopted in the above corollary. Furthermore, Corollary 3.2 enables simplifying the notation $\widetilde{\mathbb{F}}$ and $\widetilde{\mathbb{G}}$ respectively by ${\mathbb{F}}$ and ${\mathbb{G}}$ without any ambiguity. Moreover, if we define the conditional law in the enlarged probability space as $\widetilde{\mu}_{t}:=\text{Law}\big{(}\widetilde{\bm{X}}\big{|}\widetilde{\mathcal{G}}_{t}\big{)}$, then by the compatibility condition and Theorem 3.3, we have
@@
which in turn allows for using $\mu_{t}$ for $\widetilde{\mu}_{t}$ without any ambiguity.
3.4. Properties of the conditional independent copies of semimartingales
Next, we study properties of these conditional independent copies of semimartingales and explore the relation between the conditional expectation of their stochastic integrals and stochastic integrals with respect to the conditional law of semimartingales. For the sake of simplifying the notations, we will present the results in the one-dimensional case and their corresponding multi-dimensional cases can be adapted accordingly.
Theorem 3.3.
Give two ${\mathbb{R}}$-valued ${\mathbb{F}}$-semimartingales $\{X_{t}\}_{t\in[0,T]}$ and $\{Y_{t}\}_{t\in[0,T]}$ on a filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$. Let $(X^{\prime},Y^{\prime})$ and $(X^{\prime\prime},Y^{\prime\prime})$ be their conditional independent copies defined in Corollary 3.2 in the enlarged filtered probability space $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},\overline{\mathbb{P}})$ satisfying (3.6), (3.8) and (3.11). Then,
@@ 3.4. Properties of the conditional independent copies of semimartingales
    $X$, $Y$, $X^{\prime}$, $Y^{\prime}$, $X^{\prime\prime}$, $Y^{\prime\prime}$ are $\overline{\mathbb{F}}$-semimartingales.
@@
    Suppose ${\mathbb{E}}[|X_{t}|]<\infty$ and ${\mathbb{E}}[|Y_{t}|]<\infty$ for all $t\in[0,T]$, and define
    $X^{\mathcal{G}}_{t}:=\overline{\mathbb{E}}\big{[}X_{t}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{\mathbb{E}}\big{[}X_{t}\big{|}{\mathcal{G}}_{t}\big{]},\;\;Y^{\mathcal{G}}_{t}:=\overline{\mathbb{E}}\big{[}Y_{t}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{\mathbb{E}}\big{[}Y_{t}\big{|}{\mathcal{G}}_{t}\big{]}.$
@@
      If $\{X_{t}\}_{t\in[0,T]}$ is a finite variation process and ${\mathbb{E}}\big{[}\big{(}\int_{0}^{T}|{\mathrm{d}}X_{s}|\big{)}^{p}\big{]}<\infty,$ for $p\geq 1$, then $\{X^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is also a finite variation process and
@@
      $\{X^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ and $\{Y^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ are semimartingales with respect to $\overline{\mathbb{F}}$, ${\mathbb{F}}$ and ${\mathbb{G}}$.
    3. (c)
      If $\{Z_{t}\}_{t\in[0,T]}$ is an ${\mathbb{R}}$-valued ${\mathbb{F}}$-adapted càdlàg process in the filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$ and $\|X\|_{\mathcal{H}_{p}}+\|Z\|_{\mathcal{H}_{q}}<\infty$ with $\frac{1}{p}+\frac{1}{q}=1$, then
@@
The first statement in Theorem 3.3 shows that semi-martingales remain semimartingales under the enlarged filtration $\overline{F}$. (2a) and (2b) suggest that under the conditional expectation operation, the finite variation property, the semimartingale property with respect to $\overline{\mathcal{F}}$ and ${\mathcal{F}}$ are preserved and the martingale property of ${\mathcal{G}}$ is unchanged because of the compatibility condition in Assumption 2.5. In particular, (2c) and (2d) are the key equations for the derivation of the Itô’s formula for cylindrical function; they ensure crucially the equivalence between the stochastic integral of the conditional expectation of a semimartingale and the conditional expectation of the Itô’s integral of its conditional independent copy, as well as the equivalence of the quadratic variation of the conditional expectation of a semimartingale and the conditional expectation of the quadratic variation of its two conditional independent copies.
The proof of the theorem is presented in Section 5.2.
4. Proof of Theorem 2.7
Having established the properties of conditional independent copy of semimartingales, we are now ready to establish Itô’s formula for flow of conditional laws on semimratingales, through several steps. First, we will derive the formula for cylindrical functions; we will then show that cylindrical functions are dense in $C^{2,2}({\mathbb{R}}^{d}\times\mathcal{P}_{p}({\mathbb{R}}^{d}))$. Finally, appropriate localization techniques and applications of appropriate dominance convergence theorem finish the proof.
@@
where $f\in C^{2}({\mathbb{R}}^{n+d})$ and $g\in C^{2}_{b}({\mathbb{R}}^{d})$.
Next, we show that when restricted to the smaller set of functions $\Phi$, Theorem 2.7 holds. For semimartingales $\{\bm{X}_{t}\}_{t\in[0,T]}$ and $g^{(k)}\in C_{b}^{2}({\mathbb{R}}^{d})$, for $k=1,2,..,n$, by the classical Ito’s formula [37, Theorem 33, Chapter II], we have
@@
and by (2b) of Theorem 3.3, $\bm{Z}$ is an $\overline{\mathbb{F}}$-semimartingale. Using the localization argument (which is detailed in Step 2), let us assume that $\bm{Y}_{t}$ is bounded. Then applying Itô’s formula to $f\big{(}\bm{Z},\bm{Y}\big{)}$ yields
@@
Let us compute $J_{1},J_{2},J_{3}$ separately. First, we note that $g^{(i)}$ are bounded, $\bm{Y}_{t}$ is bounded, $f\in C^{2}({\mathbb{R}}^{n+d})$, we have $\nabla_{\bm{z}}f(\bm{Z}_{s-},\bm{Y}_{s-})$, $\nabla^{2}_{\bm{z}\bm{y}}f(\bm{Z}_{s-},\bm{Y}_{s-})$ and $\nabla^{2}_{\bm{z}}f(\bm{Z}_{s-},\bm{Y}_{s-})$ are bounded and we confirm the conditions for applying (2c),(2d) of Theorem 3.3. Note that for the conditional copy $\bm{X}^{\prime}$, we have

### v3

@@ 3.2. Construction of conditionally independent copies of stochastic process
and the extended $\sigma$-algebra $\overline{{\mathcal{F}}}$ by
$\overline{{\mathcal{F}}}=\sigma\{A_{0}\times A_{1}\times\cdots\times A_{n}:A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}\}.$ (3.7)
To ensure that these copies are conditionally independent, let us define the probability measure $\overline{{\mathbb{P}}}$ as follows: for $A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}$,
$\overline{{\mathbb{P}}}\big{(}A_{0}\times A_{1}\times\cdots\times A_{n}\big{)}:={\mathbb{E}}\bigg{[}\mathbbm{1}_{A_{0}}\prod_{i=1}^{n}{\mathbb{P}}(A_{i}|{\mathcal{G}})\bigg{]}.$ (3.8)
This measure is well defined since the set $\{A_{0}\times A_{1}\times\cdots\times A_{n}\}$ is a $\pi$-system. Finally, we define the completion of the above extended probability space and still use $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ to denote the completed probability space. Further, define $\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$ for the extension of $\mathcal{F}$} and $\mathcal{G}$} into this enlarged probability space as
[…]
@@
Given a probability space $(\Omega,{\mathcal{F}},{\mathbb{P}})$ and a sub $\sigma$-algebra $\mathcal{G}$$\subset{\mathcal{F}}$, there exists an enlarged completed probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ satisfying (3.6), (3.7) and (3.8); and for any random variable $R$ taking values in a standard Borel space $(E,\mathcal{E})$, there exist an $\widetilde{R}$ and $n$ conditionally independent copies $\{R^{(i)}\}_{i=1,2,..n}$ given $\widetilde{\mathcal{G}}$ such that
@@
with $\widetilde{R}$ and $\{R^{(i)}\}_{i=1,2,..n}$ satisfying the following:
  • Given $\widetilde{\mathcal{G}}$, $\widetilde{R}$ and $\big{\{}R^{(i)}\big{\}}_{i=1,2,..n}$ are conditionally independent.
  • Given $\widetilde{\mathcal{F}}$, $\widetilde{R}$ and $\big{\{}R^{(i)}\big{\}}_{i=1,2,..n}$ are conditionally independent.
  • $\big{(}R\big{|}{\mathcal{G}}\big{)}(\omega_{0})=Law$$\big{(}\widetilde{R}\big{|}\widetilde{\mathcal{G}}\big{)}(\overline{\omega})=$$\big{(}R^{(i)}\big{|}{\widetilde{\mathcal{G}}}\big{)}(\overline{\omega})=Law$$\big{(}R^{(i)}\big{|}{\widetilde{\mathcal{F}}}\big{)}(\overline{\omega})$, for all $i=1,2,..,n$, a.s.,
for $\overline{\omega}=(\omega_{0},,\omega_{1},\cdots,\omega_{n})\in\overline{\Omega}$.
Note that when $\mathcal{G}$} is a trivial $\sigma$-algebra $\{\emptyset,\Omega\}$, then the above theorem reduces to the standard construction of independent copies of stochastic processes.
3.3. Construction of independent copies of semimartingale
In the context of semimartingales, we can similarly define a corresponding version of the conditionally independent copy as follows. Take the completed filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$ with a sub-filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$ satisfying ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$, define a complete extended probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ satisfying (3.6), (3.7) and (3.8) by setting the sub $\sigma$-algebra ${\mathcal{G}}={\mathcal{G}}_{T}$. Define $\overline{{\mathbb{F}}}:=(\overline{{\mathcal{F}}}_{t})_{t\in(0,T)}$ by
$\overline{{\mathcal{F}}}_{t}:=\text{completion of }\sigma\{A_{0}\times A_{1}\times\cdots\times A_{n}:A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}_{t}\},$ (3.11)
and define $\widetilde{\mathbb{F}}:=(\widetilde{\mathcal{F}}_{t})_{t\in(0,T)}$ and $\widetilde{\mathbb{G}}:=(\widetilde{\mathcal{G}}_{t})_{t\in(0,T)}$ respectively for the extension of $\mathbb{F}$} and $\mathbb{G}$} in this enlarged probability space as
[…]
@@ 3.3. Construction of independent copies of semimartingale
[…]
@@
$(\widetilde{\bm{X}},\widetilde{\bm{Y}})(\omega_{0},\omega_{1},\cdots,\omega_{n})=(\bm{X},\bm{Y})(\omega_{0}),$
[…]
$\begin{cases}\text{Law}\big{(}\widetilde{\bm{X}},\widetilde{\bm{Y}}\big{|}\widetilde{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime},\bm{Y}^{\prime}\big{|}\widetilde{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime},\bm{Y}^{\prime}\big{|}\widetilde{\mathcal{F}}\big{)}\\ \qquad=\text{Law}\big{(}\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime}\big{|}\widetilde{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime}\big{|}\widetilde{\mathcal{F}}\big{)}\text{ a.s.,}\\ (\widetilde{\bm{X}},\widetilde{\bm{Y}}),(\bm{X}^{\prime},\bm{Y}^{\prime}),(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})\text{ are conditionally independent, given }{\mathcal{G}}\ or\ {\mathcal{F}}.\par\par\end{cases}$ (3.13)
Furthermore, if the sub $\sigma$-algebra ${\mathcal{G}}={\mathcal{G}}_{T}$, and ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}\big{|}{\mathcal{G}}_{t},$ for all $0\leq t\leq T$, then the sub-filtration $\widetilde{\mathcal{G}}\subset\widetilde{\mathcal{F}}$ satisfies $\widetilde{\mathcal{F}}_{t}\perp\!\!\!\perp\widetilde{\mathcal{G}}_{T}\big{|}\widetilde{\mathcal{G}}_{t},$ for all $0\leq t\leq T$.
[…]
@@
which in turn allows for using $\mu_{t}$ for $\widetilde{\mu}_{t}$ without any ambiguity. Therefore, from now on we sometimes will simply use the original notation to denote the natural extensions.
3.4. Examples
To get some intuition for this notion of conditionally independent copies of stochastic processes, let us see some special examples of $\mathcal{F}$} and the sub $\sigma$-algebra $\mathcal{G}$}.
Example 1. Let $A,B$ be two independent random variables with ${\mathcal{F}}=\sigma(A,B)$ and ${\mathcal{G}}=\sigma(B)$. One can enlarge the probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ to include $\{(A^{(i)},B^{(i)})\}_{i=1,2,..,n}$ satisfying (3.6), (3.7),(3.8) and (3.10). Moreover,
@@ 3.4. Examples
    $\{(A^{(i)},B^{(i)})\}_{i=1,2,..,n}$ have the same distribution as $(A,B)$.
@@
    $B^{(i)}=B$ almost surely.
  3. (3)
    $\overline{{\mathcal{F}}}=\sigma\big{(}A,B,\{A^{(i)},B^{(i)}\}_{i=1,2,..,n}\big{)}=\sigma\big{(}A,B,\{A^{(i)}\}_{i=1,2,..,n}\big{)}$.
  4. (4)
    $\{A^{(i)}\}_{i=1,2,..,n}$ and $A,B$ are independent.
Consequently, for any $X=f(A,B)$, the conditionally independent copies of $X$ are $X^{(i)}=f(A^{(i)},B^{(i)})=f(A^{(i)},B),\ \ i=1,\cdots,n.$
Example 2. Let $X$ satisfy the following stochastic differential equation
where $W,W^{(0)}$ are two independent Brownian motions, and $b$, $\sigma$, and $\sigma^{0}$ satisfy appropriate continuity conditions. Let ${\mathcal{F}}_{t}=\sigma\big{(}\{W_{s},W^{(0)}_{s}\}_{s\leq t}\big{)}$ and the sub-filtration ${\mathcal{G}}_{t}=\sigma\big{(}\{W^{(0)}_{s}\}_{s\leq t}\big{)}$, then one can enlarge the probability space to include independent Brownian motions $\big{\{}(X^{\prime},W^{\prime},{W^{(0)}}^{\prime})\big{\}}_{i=1,2,..,N}$ satisfying (3.6), (3.7),(3.8), (3.10) and (3.11). Moreover,
  1. (1)
    $X^{\prime}$ satisfies the following stochastic differential equation
    ${\mathrm{d}}X^{\prime}_{t}=b(X^{\prime}_{t}){\mathrm{d}}t+\sigma(X^{\prime}_{t}){\mathrm{d}}W^{\prime}_{t}+\sigma^{(0)}(X^{\prime}_{t}){\mathrm{d}}(W^{(0)}_{t})^{\prime}$.
  2. (2)
    $W^{(0)}=(W^{(0)})^{\prime}$ almost surely.
  3. (3)
    $\overline{{\mathcal{F}}}_{t}=\sigma\big{(}\{W_{s},W^{(0)}_{s},W^{\prime}_{s},(W^{(0}_{s})^{\prime}\}_{s\leq t}\big{)}=\sigma\big{(}\{W_{s},W^{(0}_{s},W^{\prime}_{s}\}_{s\leq t}\big{)}$
  4. (4)
    $W,W^{\prime},W^{(0)}$ are independent Brownian motions.
Consequently, $X^{\prime}$ is one conditionally independent copy of $X$ given $\mathcal{G}$}.
There are more examples of conditionally independent copies of stochastic processes in the literature of mean-field games. For instance, when defining a semimartingale by stochastic differential equations involving both idiosyncratic and common noise, conditionally independent copies are formally formulated in [14] as $\overline{\Omega}=\Omega_{0}\times\Omega_{1}\times\widetilde{\Omega}_{1}$, where the original probability space follows the structure $\Omega=\Omega_{0}\times\Omega_{1}$, and $\widetilde{\Omega}_{1}$ represents the independent copy of the probability space $\Omega_{1}$ ([14, Theorem 4.14, Volume II]).
3.5. Properties of the conditionally independent copies of semimartingales
Next, we study properties of these conditionally independent copies of semimartingales, and explore the relation between the conditional expectation of their stochastic integrals and stochastic integrals with respect to the conditional law of semimartingales. For notational simplicity, we will present the results for the one-dimensional case and their corresponding multi-dimensional cases can be obtained with appropriate adaptations.
First, by the compatibility condition and the classical results regarding filtration enlargement, we have:
Proposition 3.3.
Given an $\mathbb{R}$}-valued $\mathbb{F}$}-semimartingales $\{X_{t}\}_{t\in[0,T]}$ on a filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$. Let $X^{\prime}$ and $X^{\prime\prime}$ be its conditionally independent copies defined in Proposition 3.2 in the enlarged filtered probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}}=(\overline{{\mathcal{F}}}_{t})_{t\in[0,T]},\overline{{\mathbb{P}}})$ satisfying (3.6), (3.8) and (3.11). Then,
  1. (1)
    $X$, $^{\prime}, X′′X^{\prime\prime}$ are $\overline{{\mathbb{F}}}$-semimartingales.
  2. (2)
    Suppose ${\mathbb{E}}[|X_{t}|]<\infty$ for all $t\in[0,T]$, and define
    $X^{\mathcal{G}}_{t}:=\overline{{\mathbb{E}}}\big{[}X_{t}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{{\mathbb{E}}}\big{[}X_{t}\big{|}{\mathcal{G}}_{t}\big{]}.$
@@ 3.5. Properties of the conditionally independent copies of semimartingales
      If $\{X_{t}\}_{t\in[0,T]}$ is a finite variation process and ${\mathbb{E}}\big{[}\big{(}\int_{0}^{T}|{\mathrm{d}}X_{s}|\big{)}^{p}\big{]}<\infty$ for $p\geq 1$, then $\{X^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is also a finite variation process, and
@@
      $\{X^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is a semimartingale with respect to $\overline{{\mathbb{F}}}$, $\mathbb{F}$}, and $\mathbb{G}$}.
Remark 3.4.
Proposition 3.3 shows that semimartingales remain semimartingales under the enlarged filtration $\overline{{\mathbb{F}}}$. This is closely related to the so called $H^{\prime}$-hypothesis where any $\mathbb{F}$}-martingale is an $\overline{{\mathbb{F}}}$-semimartingale with the enlargement filtration from $\mathbb{F}$} to $\overline{{\mathbb{F}}}$ (see for instance [3, 29]). Here we have obtained a slightly stronger result that any $\mathbb{F}$}-semimartingale $X$ and its conditionally independent copy remain $\overline{{\mathbb{F}}}$-semimartingales; meanwhile the $\mathbb{G}$}-semimartingale $X^{\mathcal{G}}$ (the projection taken with respect to the $\sigma$-algebra $\mathcal{G}$} as in [8]) remains $\overline{{\mathbb{F}}}$-semimartingale. Despite the apparent similarity, we are unable to locate exactly the same statement, and will provide the proof for completeness.
Theorem 3.5.
Give two $\mathbb{R}$}-valued $\mathbb{F}$}-semimartingales $\{X_{t}\}_{t\in[0,T]}$ and $\{Y_{t}\}_{t\in[0,T]}$ on a filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$. Let $(X^{\prime},Y^{\prime})$ and $(X^{\prime\prime},Y^{\prime\prime})$ be their conditionally independent copies defined in Proposition 3.2 in the enlarged filtered probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}}=({\mathcal{F}}_{t})_{t\in[0,T]},\overline{{\mathbb{P}}})$ satisfying (3.6), (3.8) and (3.11). Suppose ${\mathbb{E}}[|X_{t}|]<\infty$ and ${\mathbb{E}}[|Y_{t}|]<\infty$ for all $t\in[0,T]$.
  1. (1)
    If $\{Z_{t}\}_{t\in[0,T]}$ is an $\mathbb{R}$}-valued $\mathbb{F}$}-adapted càdlàg process in the filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$, and $\|X\|_{\mathcal{H}_{p}}+\|Z\|_{\mathcal{H}_{q}}<\infty$ with $\frac{1}{p}+\frac{1}{q}=1$, then
@@
Remark 3.6.
Statements (1) and (2) in Theorem 3.5 are the key equations for deriving the Itô’s formula (Theorem 2.11. They ensure crucially a) the equivalence between the stochastic integral of the conditional expectation of a semimartingale and the conditional expectation of the Itô’s integral of its conditionally independent copy, and b) the equivalence between the quadratic variation of the conditional expectation of a semimartingale and the conditional expectation of the quadratic variation of its two conditionally independent copies.
The proof of Theorem 3.5 is presented in Section 5.3.
4. Proof of Main Result (Theorem 2.11)
Having established necessary properties of conditionally independent copies of semimartingales, we are now ready to prove Theorem 2.11, the Itô’s formula for flow of conditional laws on semimartingales, through several steps. First, we will derive the formula for cylindrical functions; we will then show that cylindrical functions are dense in $C^{2,2}({\mathbb{R}}^{d}\times\mathcal{P}_{p}({\mathbb{R}}^{d}))$; finally, with appropriate localization technique and dominated convergence theorem, we will finish the proof.
@@
where $f\in C^{2}({\mathbb{R}}^{n+d})$ and $g=(g^{(1)},..,g^{(n)})$, with $g^{(i)}\in C^{2}_{b}({\mathbb{R}}^{d})$.
We will first show that for smooth cylindrical functions $\Phi$, Theorem 2.11 holds. Indeed, for semimartingales $\{\bm{X}_{t}\}_{t\in[0,T]}$ and $g^{(i)}\in C_{b}^{2}({\mathbb{R}}^{d})$ with $i=1,2,..,n$, by the classical Itô’s formula [51, Theorem 33, Chapter II],
@@
then by (2b) of Theorem 3.5, $\bm{Z}$ is an $\overline{{\mathbb{F}}}$-semimartingale.
Let us for the moment assume that $\bm{Y}_{t}$ is bounded. Then applying Itô’s formula to $f\big{(}\bm{Z},\bm{Y}\big{)}$ yields
@@
We will compute $J_{1},J_{2},$ and $J_{3}$ separately. First, note that both $g^{(i)}$ and $\bm{Y}_{t}$ are bounded, and $f\in C^{2}({\mathbb{R}}^{n+d})$, therefore $\nabla_{\bm{z}}f(\bm{Z}_{s-},\bm{Y}_{s-})$, $\nabla^{2}_{\bm{z}\bm{y}}f(\bm{Z}_{s-},\bm{Y}_{s-})$, and $\nabla^{2}_{\bm{z}}f(\bm{Z}_{s-},\bm{Y}_{s-})$ are bounded, and the conditions for applying (1) and (2) of Theorem 3.5 are satisfied.
For the conditionally independent copy $\bm{X}^{\prime}$, we have